OpenStudy (anonymous):

Multiply and simplify: 9^4*9^-13

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 9^4 \times 9^{-13}= 9^{~4+(-13)}=9^{~4-13}=9^{-9} }\)

OpenStudy (solomonzelman):

Then there is a rule, that says, \(\LARGE\color{blue}{ a^{-b}=\frac{1}{a^b}}\) Knowing the above rule, \(\LARGE\color{blue}{ 9^{-9}=\frac{1}{9^9}}\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ 9^{9}=387420489}\) and knowing this, you will get 1 / 387420489

OpenStudy (solomonzelman):

this is a very small number that is not so far from ZERO

OpenStudy (solomonzelman):

Anything else I can help you with ?

OpenStudy (anonymous):

That is so wonderful! Thank you! This is super to understand now! I appreciate the help

OpenStudy (solomonzelman):

Anytime! This is what I am here for (and I am here for to practice my maths, so thank YOU! )